OpenStudy (anonymous):

The youth group is going on a trip to the state fair. The trip costs $62. Included in that price is $14 for a concert ticket and the cost of 2 passes, one for the rides and one for the game booths. Each of the passes costs the same price. Write an equation representing the cost of the trip, and determine the price of one pass. Solve your equation by showing your work and steps

OpenStudy (study312):

62 - 14 = 48 48 ÷ 2 = 24. it costs $24

OpenStudy (study312):

I mean 14 + 2x = 62

OpenStudy (anonymous):

that's all?

OpenStudy (study312):

well from what you got from my answer try to make a equation and solve it

OpenStudy (anonymous):

which 1 u had like 3 answers

OpenStudy (study312):

like this 14 + 2x = 62 solve it.

OpenStudy (study312):

because its asking you to make an equation and solve it

OpenStudy (study312):

now if you don't like 14 + 2x = 62 you can try and make another equation to find the value of one pass

OpenStudy (anonymous):

isn't it asking me to make an equation about the problem its asking me?

OpenStudy (study312):

yes, isn't that what I just said?

OpenStudy (anonymous):

oh yea lol

OpenStudy (anonymous):

how do Ik wat x is?

OpenStudy (study312):

so do you get it? make a equation from the question and solve it, I showed you the answer so that should help make the equation and solve it

OpenStudy (study312):

well like this 14 + 2x = 62 14 - 14 + 2x = 62 - 14. subtract 14 from both sides 2x = 48 48 ÷ 2 = 24. x = 24

OpenStudy (study312):

so your answer is what I just commented

OpenStudy (anonymous):

oh ty I appreciate it

OpenStudy (study312):

No problem